# MnLargeSymbols’164 MnLargeSymbols’171

### v1

@@ 3.3 Construction by decomposition
where the first terms ${\bm{A}}_{i}^{b}$ lift some suitable boundary condition into the domain, and the second terms ${\bm{A}}_{i}^{0}$ are correction terms with homogeneous boundary conditions.
Specifically, the ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ are piecewise smooth fields (in the sense of Section 3.1) which lift the following boundary conditions:
${\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0~~\text{ on }\partial\Omega\qquad\text{ and }\qquad\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}.$ (38)
Notice that the first condition may be reformulated using a surface operator (see, e.g., [6, Cor. 3.1.16]) so that it only involves boundary values of ${\bm{A}}^{b}_{i}$ indeed. We will say that the ${\bm{A}}_{i}^{b}$ are lifted boundary potentials.
@@
If (38) and (39) hold, then the potentials ${\bm{A}}_{i}$ satisfy $\operatorname{{\bf curl}}{\bm{A}}_{i}\in\mathfrak{H}^{1}$ and
@@
The equality $\operatorname{div}{\bm{w}}_{i}=0$ is obvious, and $\operatorname{{\bf curl}}{\bm{w}}_{i}=0$ is clear from (39). The boundary condition ${\bm{w}}_{i}\in H_{0}(\operatorname{div};\Omega)$ is also easily verified, indeed it is satisfied by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}$ thanks to (38), and by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{0}$ thanks to the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$. This shows that the fields ${\bm{w}}_{i}$ are indeed tangent-harmonic. To show the relations (40) we first consider the lifted boundary part. Since it is piecewise smooth, Stokes’ formula (32) yields
@@
where the last equality is (38). The desired relation (40) then follows from the fact that the flux functionals $\Phi_{j}$ vanish for the ${\bm{A}}^{0}_{i}$ parts, thanks to (35). The linear independence of the fields ${\bm{w}}_{i}$’s follows from (40), and the fact that they span $\mathfrak{H}^{1}$ follows from (10). ∎
@@ 3.4 Lifted boundary potentials
where $\hat{{\bm{n}}}=-\hat{{\bm{e}}}_{r}$ is the unit normal vector pointing outside of $\mathcal{O}$ on its $\{r=0\}$ boundary, and
$\int_{-1}^{1}\hat{\bm{\tau}}\cdot\widehat{{\bm{A}}}^{b}(0,0,z)\ \mathrm{d}z=1$ (42)
where $\hat{\bm{\tau}}=\hat{{\bm{e}}}_{z}$ is the unit vector tangent to the $\{r=0,\theta=0\}$ curve. Since $\widehat{{\bm{A}}}^{b}$ is compactly supported, it vanishes on every boundary of $\mathcal{O}$ with the exception of the surface $\{r=0\}$ which is contained in $\mathcal{O}$: it is on this boundary that the above boundary conditions are imposed.
@@
where $\mathcal{F}^{1}_{i}(\hat{\bm{u}}):=((DT_{i})^{-T}\hat{\bm{u}})\circ T_{i}^{-1}$ is the covariant Piola transform (with $DT_{i}$ the Jacobian matrix of $T_{i}$) corresponding to the pushforward operator for 1-form proxy fields [22]. Let us denote by $\mathcal{O}_{i,a}$, $a=1,\dots,m_{i}$ the subdomains of $\mathcal{O}$ on which $T_{i}$ is smooth. On the corresponding domains $X_{i,a}=T_{i}(\mathcal{O}_{i,a})$ the field ${\bm{A}}_{i}^{b}$ is smooth as well.
@@
The field ${\bm{A}}_{i}^{b}$ just defined satisfies the properties stated above, namely it is $H(\operatorname{{\bf curl}};\Omega)$ and the boundary conditions (38) hold.
@@
To verify that ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ we need to show that (i) the tangential trace of the pushforward $\mathcal{F}_{i}^{1}(\widehat{{\bm{A}}}^{b})({\bm{x}})$ on $\partial X_{i}\setminus\partial\Omega$ vanishes, and (ii) on every interface $f=\overline{X_{i,a}}\cap\overline{X_{i,b}}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, we have
@@
The former property is clear since the logical field (43) vanishes on a neighborhood of $\partial\mathcal{O}\setminus\{r=0\}$, and the latter property follows from from the fact that a tangent vector on $f$ is of the form
@@
where $\hat{\bm{\tau}}$ is a vector tangent to the logical surface $\hat{f}=T_{i}^{-1}(f)$, and where the Jacobian matrix of $T_{i}$ can be evaluated from both subdomains $\mathcal{O}_{i,a}$ or $\mathcal{O}_{i,b}$ since the directional derivative is taken along their common interface. The continuity of the tangential traces on $f$ then follows from the definition of the pushforward.
Finally to verify the boundary conditions we use the commutation relation
@@
satisfied by the 2-form pushforward operator $\mathcal{F}^{2}_{i}$, see e.g. [22]. The relations (38) then follow from (41), (42), and the fact that line and surface integrals of smooth vector fields are respectively preserved by 1-form and 2-form pushforward operators [24, Ch. 16.6]. ∎
@@ 3.5 Correction potentials
The system (46) defines a unique field ${\bm{A}}_{i}^{0}$ which satisfies (39).
@@
The well-posedness follows by applying [5, Theorem 3.2], to the adjoint (homogeneous) de Rham complex (2), which is possible thanks to the properties stated in [5, Sec. 4.2]. By taking respectively ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{i}$ and ${\bm{v}}={\bm{p}}_{i}\in\mathfrak{H}^{2}$, we then observe that the solution satisfies $\sigma_{i}=0$ and ${\bm{p}}_{i}=0$. The relation (39) readily follows. ∎
Combining Propositions 3.4, 3.5 and 3.6 yields the following result.
@@
The fields ${\bm{A}}_{i}={\bm{A}}^{b}_{i}+{\bm{A}}^{0}_{i}$ obtained with the above construction are vector potentials for the tangent harmonic fields. In particular, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$, $i=1,\dots,\beta_{1}$, form a basis of $\mathfrak{H}^{1}$ that is dual to the flux functionals $\Phi_{i}$.
@@ 4 Structure-preserving discrete harmonic potentials
We end this article by describing how the above construction can be implemented in a discrete setting. The approach proposed here applies to structure-preserving finite element spaces that share a certain number of properties listed in Sections 4.1 and 4.3 below.
Assuming that the domain $\Omega$ can be triangulated with a simplicial mesh, these properties hold for instance in the case of Whitney [36, 9] and Nédélec [29] finite element spaces, and more generally for high order FEEC spaces [4], but they are not restricted to these cases.
4.1 Discretization of the homogeneous (dual) and extended dual complexes
@@
Notice that up to a change of signs in the grad and div operators, the latter coincides with the primal complex (1). For simplicity we may disregard these sign changes and consider the plain differential operators below. However, we should not confuse the primal and the extended dual complexes.
We then consider a discretization of the extended dual complex equipped with degrees of freedom that are commuting and compatible with the Dirichlet boundary conditions. By this we mean essentially two properties. First, that we are given a discrete subcomplex of (48),
$0\xleftarrow{~~}\tilde{V}^{3}_{h}\xleftarrow{~\operatorname{div}~}\tilde{V}^{2}_{h}\xleftarrow{~\operatorname{{\bf curl}}~}\tilde{V}^{1}_{h}\xleftarrow{~\operatorname{{\bf grad}}~}\tilde{V}^{0}_{h}\xleftarrow{~~}0$ (49)
with spaces of piecewise continuous functions $\tilde{V}^{\ell}_{h}\subset C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ (associated with some mesh $\mathcal{M}_{h}$ of possibly curved or polygonal Lipschitz cells, as described in Section 3.1) equipped with linear degrees of freedom
$\tilde{\sigma}^{\ell}_{i}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\mathbb{R},\qquad 1\leq i\leq N_{\ell}:=\dim(\tilde{V}^{\ell}_{h}),$ (50)
satisfying the following commuting relation: for all $v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ with $\tilde{d}^{\ell}v\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell+1}$, it holds
$\tilde{\sigma}^{\ell+1}_{i}(\tilde{d}^{\ell}v)=\sum_{j=1}^{N_{\ell}}\tilde{\mathbb{D}}^{\ell}_{ij}\tilde{\sigma}^{\ell}_{j}(v),\qquad 1\leq i\leq N_{\ell+1},$ (51)
@@
with a fixed $N_{\ell+1}\times N_{\ell}$ matrix $\mathbb{D}^{\ell}$ that represents the differential operator at the discrete level. We note that a relation of the form (51) actually implies that the corresponding projection operators $\tilde{\Pi}^{\ell}:(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\tilde{V}^{\ell}_{h}$, defined by
$\tilde{\sigma}^{\ell}_{i}(\tilde{\Pi}^{\ell}v)=\tilde{\sigma}^{\ell}_{i}(v),\qquad 1\leq i\leq N_{\ell},$ (52)
@@
commute with the differential operators. The second property is that the degrees of freedom lead to projection operators that preserve the homogenous boundary conditions, in the sense that we have
$\tilde{\Pi}^{\ell}(C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}_{0})=(\tilde{V}^{\ell}_{h}\cap\tilde{V}^{\ell}_{0})=:\tilde{V}^{\ell}_{h,0}.$ (53)
@@
In particular, the same projection operators yield two commuting diagrams: one for the dual complex (with homogeneous boundary conditions)
[…]
and another one for the extended dual complex,
[…]
4.2 Flux functionals on discrete surfaces
To study the linear independence of our discrete vector potentials, we assume in addition that there exist surfaces $\Sigma_{h,i}$ made of smooth surface with corners, with the following properties: (i) the line integrals over the boundaries $\partial\Sigma_{h,i}$ can be expressed as linear combinations of degrees of freedom in $\tilde{V}^{1}_{h}$, (ii) these boundaries are included in the domain boundary ($\partial\Sigma_{h,i}\subset\partial\Omega$) and (iii) they coincide with those of the smooth reciprocal surfaces $\Sigma_{i}$ on the parametrized domains $X_{j}$, in the sense that
$X_{j}\cap\partial\Sigma_{h,i}=X_{j}\cap\partial\Sigma_{i},\qquad 1\leq i,j\leq\beta_{1}.$ (56)
We remind that these intersections are empty for $i\neq j$, see (31).
In the case of Whitney, Nédélec or more general FEEC finite elements spaces, the usual degrees of freedom involve line integrals along edges of the mesh, so that the first property holds for surfaces $\Sigma_{h,i}$ made of faces of the mesh $\mathcal{M}_{h}$. The second property poses no particular difficulty, and the last one can then be fulfilled by choosing a parametrization of $\Gamma_{i}$ such that the smooth curves $X_{i}\cap\partial\Sigma_{i}$ are included in a single edge of the boundary $\partial\Sigma_{h,i}$.
In turn, this implies that the finite element projection $\tilde{\Pi}^{1}$ interpolates the corresponding line integrals: indeed one has
$\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}(\tilde{\Pi}^{1}{\bm{a}})=\sum_{j=1}^{N_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}({\bm{a}})=\int_{\partial\Sigma_{h,i}}{\bm{\tau}}\cdot{\bm{a}}$
@@ 4.2 Flux functionals on discrete surfaces
for all ${\bm{a}}\in C^{0}(\mathcal{M}_{h})\cap\tilde{V}^{1}$, where the coefficients $c_{i,j}$ correspond to the decomposition of the line integrals in terms of degrees of freedom. Moreover, applying Stokes’ theorem on each smooth face $f\in F(\Sigma_{h,i})$, we compute

### v2

@@ 3.3 Construction by decomposition
where the first terms lift some suitable boundary condition into the domain, and the second are correction terms with homogeneous boundary conditions.
Specifically, the ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$ are piecewise smooth fields which lift the following boundary conditions:
${\bm{n}}_{\partial\Omega}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0~~\text{ on }\partial\Omega\qquad\text{ and }\qquad\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\delta_{i,j},\qquad 1\leq i,j\leq\beta_{1}$ (46)
where ${\bm{n}}_{\partial\Omega}$ is the unit normal vector on the Lipschitz surface $\partial\Omega$, pointing outside of $\Omega$. Notice that the first condition may be reformulated using a surface operator (see, e.g., [8, Cor. 3.1.16]) so that it only involves boundary values of ${\bm{A}}^{b}_{i}$ indeed. We will say that the ${\bm{A}}_{i}^{b}$ are lifted boundary potentials.
@@
If (46) and (47) hold, then the potentials (45) satisfy $\operatorname{{\bf curl}}{\bm{A}}_{i}\in\mathfrak{H}^{1}$ and
@@
The equality $\operatorname{div}{\bm{w}}_{i}=0$ is obvious, and $\operatorname{{\bf curl}}{\bm{w}}_{i}=0$ is clear from (47). The boundary condition ${\bm{w}}_{i}\in H_{0}(\operatorname{div};\Omega)$ is also easily verified, indeed it is satisfied by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}$ thanks to (46), and by $\operatorname{{\bf curl}}{\bm{A}}_{i}^{0}$ thanks to the inclusion $\operatorname{{\bf curl}}H_{0}(\operatorname{{\bf curl}};\Omega)\subset H_{0}(\operatorname{div};\Omega)$. This shows that the fields ${\bm{w}}_{i}$ are indeed tangent and harmonic. To show the relations (48) we first consider the lifted boundary part. Since it is piecewise smooth over the mesh $\mathcal{X}_{i}$, Stokes’ formula (38) yields
@@
where the last equality is (46). The desired relation (48) then follows from the fact that the flux functionals $\Phi_{j}$ vanish for the ${\bm{A}}^{0}_{i}$ parts thanks to (41). The linear independence of the fields ${\bm{w}}_{i}$ readily follows, and the fact that they span $\mathfrak{H}^{1}$ is a consequence of de Rham’s Theorem 2.1. ∎
@@ 3.4 Lifted boundary potentials
where $\hat{{\bm{n}}}=-\hat{{\bm{e}}}_{r}$ is the unit normal vector pointing outside of $\mathcal{O}$ on its $\{r=0\}$ boundary (see Figure 2), and
$\int_{-1}^{1}\hat{{\bm{e}}}_{z}\cdot\widehat{{\bm{A}}}^{b}(0,\theta,z)\ \mathrm{d}z=1,\qquad\forall\theta\in\mathbb{T}.$ (50)
Note that since $\widehat{{\bm{A}}}^{b}$ is compactly supported in $\mathcal{O}$, it vanishes on every face with the exception of the face $\{r=0\}$ which is contained in $\mathcal{O}$: it is on this face that the above boundary conditions are imposed.
@@
where $\mathcal{F}^{1}_{i}(\hat{\bm{u}}):=((DT_{i})^{-T}\hat{\bm{u}})\circ T_{i}^{-1}$ is the covariant Piola transform (with $DT_{i}$ the Jacobian matrix of $T_{i}$) corresponding to the pushforward operator for 1-form proxy fields [28].
@@
The field ${\bm{A}}_{i}^{b}$ just defined satisfies the properties stated above: it is piecewise $C^{\infty}$ over the mesh $\mathcal{X}_{i}$ in (33), it belongs to $H(\operatorname{{\bf curl}};\Omega)$ and the boundary conditions (46) hold.
@@
By construction ${\bm{A}}_{i}^{b}$ vanishes on $\overline{X}_{i,0}$, and its smoothness on $\overline{X}_{i,a}$, $a=1,\dots m_{i}$, follows from the smoothness of the logical potential (51) and that of the mapping $T_{i}$ and its inverse on the subdomains $\overline{\mathcal{O}}_{i,a}$ and $\overline{X}_{i,a}=T_{i}(\overline{\mathcal{O}}_{i,a})$. To verify that ${\bm{A}}_{i}^{b}\in H(\operatorname{{\bf curl}};\Omega)$, according to [8, Prop. 2.2.32] we need to show that (i) the tangential trace of the potential (52) vanishes on $\partial X_{i}\setminus\partial\Omega$, and (ii) on every interface $f=\overline{X}_{i,a}\cap\overline{X}_{i,b}$ between two subdomains, and every vector ${\bm{\tau}}_{f}$ tangent to $f$, we have
@@
The former property is clear since the field (51) vanishes on a neighborhood of $\partial\mathcal{O}\setminus\{r=0\}$, and the latter property follows from from the fact that a tangent vector on $f$ is of the form
@@
where $\hat{\bm{\tau}}$ is a vector tangent to the surface $\hat{f}=T_{i}^{-1}(f)$, and where the Jacobian matrix of $T_{i}$ can be evaluated from both subdomains $\mathcal{O}_{i,a}$ or $\mathcal{O}_{i,b}$ since the directional derivative is taken along their common interface. The continuity of the tangential traces on $f$ then follows from the definition of the pushforward.
We finally turn to the boundary conditions (46). Using the commutation relation
@@
satisfied by the 2-form pushforward operator $\mathcal{F}^{2}_{i}$, see e.g. [28], and the fact that surface integrals of smooth vector fields are preserved by 2-form pushforwards [30, Ch. 16.6], we verify that the first boundary condition ${\bm{n}}_{\partial\Omega}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}^{b}=0$ is satisfied as a consequence of its parametric counterpart (49). To verify the second one, we remind that the intersection of $\partial\Sigma_{j}$ with $X_{i}$ decomposes into $n(i,j)$ disjoint curves according to (35)–(37) (or (34) in the simpler case where $n(i,j)=\delta_{i,j}$) that is,
$X_{i}\cap\partial\Sigma_{j}=\cup_{1\leq p\leq n(i,j)}\tilde{\Gamma}(i,j,p)\quad\text{ with }~\tilde{\Gamma}(i,j,p)=T_{i}\big(\{0\}\times\{\theta(i,j,p)\}\times(-1,1)\big)$
each curve being oriented by a unit vector induced by the global orientation of $\Sigma_{j}$ as defined by its normal vector field ${\bm{n}}_{\Sigma_{j}}$. Using the fact that line integrals are preserved by 1-form pushforwards, we then write
$\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}\int_{\tilde{\Gamma}(i,j,p)}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}\int_{-1}^{1}\hat{\bm{\tau}}(i,j,p)\cdot\widehat{{\bm{A}}_{i}^{b}}(0,\theta(i,j,p),z)\ \mathrm{d}z$
where $\hat{\bm{\tau}}(i,j,p)$ is the tangent vector of the straight curve $T_{i}^{-1}(\tilde{\Gamma}(i,j,p))$ in the direction induced by that of $\tilde{\Gamma}(i,j,p)$, that is,
$\hat{\bm{\tau}}(i,j,p)={\rm ind}(i,j,p)\hat{{\bm{e}}}_{z}$
according to (36). Using (50) and (37) then yields the desired result:
$\int_{\partial\Sigma_{j}}{\bm{\tau}}\cdot{\bm{A}}_{i}^{b}=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)\int_{-1}^{1}\hat{{\bm{e}}}_{z}\cdot\widehat{{\bm{A}}_{i}^{b}}(0,\theta(i,j,p),z)\ \mathrm{d}z=\sum_{p=1}^{n(i,j)}{\rm ind}(i,j,p)=\delta_{i,j}.$
@@ 3.5 Correction potentials
System (54) defines a unique field ${\bm{A}}_{i}^{0}$ which satisfies (47).
@@
The well-posedness follows by applying [7, Theorem 3.2], to the adjoint (homogeneous) de Rham complex (2), which is possible thanks to the properties stated in [7, Sec. 4.2]. By taking respectively ${\bm{v}}=\operatorname{{\bf grad}}\sigma_{i}$ and ${\bm{v}}={\bm{p}}_{i}\in\mathfrak{H}^{2}$, we then observe that the solution satisfies $\sigma_{i}=0$ and ${\bm{p}}_{i}=0$. The relation (47) readily follows. ∎
3.6 Main result and invariance principles
By combining Propositions 3.8, 3.9 and 3.10, which hold under Assumption 3.3, we have proven our main result.
@@ 3.6 Main result and invariance principles
The fields ${\bm{A}}_{i}={\bm{A}}^{b}_{i}+{\bm{A}}^{0}_{i}$ constructed over the tunnel curves $\Gamma_{i}$, $i=1,\dots\beta_{1}$ as described above, are free vector potentials for the tangent harmonic fields. More precisely, the fields ${\bm{w}}_{i}:=\operatorname{{\bf curl}}{\bm{A}}_{i}$ form a basis for the space $\mathfrak{H}^{1}$ which is dual to the flux functionals $\Phi_{i}$ associated with the reciprocal surfaces, in the sense that
$\Phi_{j}(\operatorname{{\bf curl}}{\bm{A}}_{i})=\int_{\Sigma_{j}}{\bm{n}}\cdot\operatorname{{\bf curl}}{\bm{A}}_{i}=\delta_{i,j}$ (55)
holds for all $1\leq i,j\leq\beta_{1}$.
Our construction satisfies two (dual) invariance principles. The first one is with respect to the tunnel curves: if $\Gamma_{i}$ and $\Gamma^{\prime}_{i}$, $i=1,\dots\beta_{1}$, are two collections of piecewise smooth closed boundary curves satisfying the properties of Assumption 3.3 with the same reciprocal surfaces $\Sigma_{i}$, then their associated harmonic potentials are a priori different but their harmonic fields coincide
$\operatorname{{\bf curl}}{\bm{A}}_{i}=\operatorname{{\bf curl}}{\bm{A}}^{\prime}_{i},\quad i=1,\dots\beta_{1},$ (56)
since it follows from Theorem 3.1 that the fluxes $\Phi_{i}$ are unisolvent on the harmonic space $\mathfrak{H}^{1}$.
This invariance principle actually holds with respect to the 1-chain homology classes: indeed if $\Gamma_{i}$ and $\Gamma^{\prime}_{i}$ are two collections of piecewise smooth closed boundary curves that represent the same basis $\Gamma_{i}]$ of the first homology group $\mathcal{H}_{1}(\overline{\Omega})$, then we may consider the (unique) basis $\Sigma_{i}]$ of the relative homology group $\mathcal{H}_{2}(\overline{\Omega},\partial\Omega)$ which is dual in the sense of the intersection pairing (29), and any collection of piecewise smooth surfaces $\Sigma_{i}$ which represent this dual basis while intersecting nicely all the tunnel curves. Because the intersection pairing coincides with the algebraic intersection number of any pair of representative cycles, we find that the surfaces $\Sigma_{i}$ are reciprocal to both curve collections, so that the associated harmonic fields coincide again.
A dual invariance principle holds with respect to the reciprocal surfaces: indeed if $\Sigma_{i}$ and $\Sigma^{\prime}_{i}$, $i=1,\dots\beta_{1}$, are two collections of piecewise smooth surfaces wich are reciprocal to the same tunnel curves $\Gamma_{i}$ in the sense of Assumption 3.3, then the harmonic fields ${\bm{w}}_{i}=\operatorname{{\bf curl}}{\bm{A}}_{i}$ associated with these curves have the same fluxes across both surface collections:
$\int_{\Sigma_{i}}{\bm{n}}\cdot{\bm{w}}_{i}=\int_{\Sigma^{\prime}_{i}}{\bm{n}}\cdot{\bm{w}}_{i},\quad i=1,\dots\beta_{1},$ (57)
as an immediate consequence of Theorem 3.1.
Again, it follows from the properties of the intersection pairing (29) that this dual invariance principle actually holds with respect to the homology classes of 2-chains modulo the boundary.
@@ 4 Structure-preserving discrete harmonic potentials
We end this article by describing how the above construction can be implemented in a discrete setting. The approach proposed here applies to structure-preserving finite element spaces that share a certain number of properties listed in Sections 4.1 and 4.2 below.
Assuming that the domain $\Omega$ can be triangulated with a simplicial mesh, these properties hold for instance in the case of Whitney [44, 12] and Nédélec [37] finite element spaces, and more generally high order FEEC spaces [6], but they are not restricted to these cases.
4.1 Discretization of the homogeneous and extended dual complexes
@@
Notice that up to a change of signs in the grad and div operators, the latter coincides with the primal complex (1). For simplicity we may disregard these sign changes and consider the plain differential operators below. We distinguish, however, the primal and the extended dual complexes which play different roles.
We then consider a discrete subcomplex of (59),
$0\xleftarrow{~~}\tilde{V}^{3}_{h}\xleftarrow{~\operatorname{div}~}\tilde{V}^{2}_{h}\xleftarrow{~\operatorname{{\bf curl}}~}\tilde{V}^{1}_{h}\xleftarrow{~\operatorname{{\bf grad}}~}\tilde{V}^{0}_{h}\xleftarrow{~~}0,$ (60)
made of finite-dimensional spaces of piecewise $C^{\infty}$ functions over some mesh $\mathcal{M}_{h}$ of $\Omega$, made of cells with Lipschitz and piecewise smooth boundaries as in Definition 3.1. Thus, we consider discrete spaces
$\tilde{V}^{\ell}_{h}\subset C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell},\qquad 0\leq\ell\leq 3,$ (61)
with the following properties:
  • (P1)
    $\mathcal{M}_{h}$ is finer than every mesh $\mathcal{X}_{i}$ from (33), in the sense that for all $1\leq i\leq\beta_{i}$ it holds
    $C^{\infty}(\mathcal{X}_{i})\subseteq C^{\infty}(\mathcal{M}_{h})$ (62)
    moreover $\mathcal{M}_{h}$ is compatible with every surface $\Sigma_{i}$ in the sense of Definition 3.2,
  • (P2)
    each space $\tilde{V}^{\ell}_{h}$ is equipped with degrees of freedom
    $\tilde{\sigma}^{\ell}_{i}:(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\mathbb{R},\qquad 1\leq i\leq\tilde{N}_{\ell}:=\dim(\tilde{V}^{\ell}_{h}),$ (63)
    satisfying relations of the form
    $\tilde{\sigma}^{\ell+1}_{i}(\tilde{d}^{\ell}v)=\sum_{j=1}^{\tilde{N}_{\ell}}\tilde{\mathbb{D}}^{\ell}_{ij}\tilde{\sigma}^{\ell}_{j}(v),\qquad 1\leq i\leq\tilde{N}_{\ell+1},$ (64)
    for all $v\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}$ with $\tilde{d}^{\ell}v\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell+1}$, and some $\tilde{N}_{\ell+1}\times\tilde{N}_{\ell}$ matrices $\tilde{\mathbb{D}}^{\ell}$.
@@
The matrices $\tilde{\mathbb{D}}^{\ell}$ represent the differential operators $\tilde{d}^{\ell}$ at the discrete level, and we note that the relations (64) imply that the corresponding projection operators $\tilde{\Pi}^{\ell}:(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell})\to\tilde{V}^{\ell}_{h}$, defined by
$\tilde{\sigma}^{\ell}_{i}(\tilde{\Pi}^{\ell}v)=\tilde{\sigma}^{\ell}_{i}(v),\qquad 1\leq i\leq\tilde{N}_{\ell}$, (65)
commute with the differential operators. Further, we assume that
  • (P3)
@@
    the projection operators preserve the homogenous boundary conditions, in the sense that
    $\tilde{\Pi}^{\ell}(C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{\ell}_{0})=\tilde{V}^{\ell}_{h}\cap\tilde{V}^{\ell}_{0}=:\tilde{V}^{\ell}_{h,0}$ (66)
  • (P4)
    the line integrals over the reciprocal surface boundaries $\partial\Sigma_{i}$, $1\leq i\leq\beta_{1}$, can be expressed as linear combinations of degrees of freedom in $\tilde{V}^{1}_{h}$.
We observe that in practice Property (P1) is not very restrictive since in most cases one can subdivide the surfaces $\Sigma_{i}$ into smaller pieces to make them compatible with a fine mesh. In the case of Whitney, Nédélec or more general FEEC finite elements, the usual degrees of freedom satisfy property (P2) and (P3). Moreover the curl-conforming degrees of freedom $\tilde{\sigma}^{1}_{i}$ involve line integrals along edges of the mesh, so that property (P4) holds as long as one can build the reciprocal surfaces $\Sigma_{i}$ by assembling some faces of the mesh $\mathcal{M}_{h}$ (which readily makes them compatible).
@@
It follows from properties (P2) and (P3) that the same projection operators yield two commuting diagrams: one for the dual complex (58) (with homogeneous boundary conditions)
$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{3}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{2}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}_{0}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{0}_{0}~\phantom{\big|}$$\tilde{V}^{3}_{h,0}~\phantom{\big|}$$\tilde{V}^{2}_{h,0}~\phantom{\big|}$$\tilde{V}^{1}_{h,0}~\phantom{\big|}$$\tilde{V}^{0}_{h,0}~\phantom{\big|}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\tilde{\Pi}^{3}$$\tilde{\Pi}^{2}$$\tilde{\Pi}^{1}$$\tilde{\Pi}^{0}$ (67)
and another one for the extended dual complex (59),
$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{3}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{2}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1}~\phantom{\big|}$$C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{0}~\phantom{\big|}$$\tilde{V}^{3}_{h}~\phantom{\big|}$$\tilde{V}^{2}_{h}~\phantom{\big|}$$\tilde{V}^{1}_{h}~\phantom{\big|}$$\tilde{V}^{0}_{h}~.\phantom{\big|}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\operatorname{div}$$\operatorname{{\bf curl}}$$\operatorname{{\bf grad}}$$\tilde{\Pi}^{3}$$\tilde{\Pi}^{2}$$\tilde{\Pi}^{1}$$\tilde{\Pi}^{0}$ (68)
Our assumptions also imply that the projection $\tilde{\Pi}^{1}$ interpolates the line integrals along the boundaries $\partial\Sigma_{i}$, $1\leq i\leq\beta_{1}$: one has indeed
$\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot\tilde{\Pi}^{1}{\bm{a}}=\sum_{j=1}^{\tilde{N}_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}(\tilde{\Pi}^{1}{\bm{a}})=\sum_{j=1}^{\tilde{N}_{1}}c_{i,j}\tilde{\sigma}^{1}_{j}({\bm{a}})=\int_{\partial\Sigma_{i}}{\bm{\tau}}\cdot{\bm{a}},\qquad{\bm{a}}\in C^{\infty}(\mathcal{M}_{h})\cap\tilde{V}^{1},$
@@
where the coefficients $c_{i,j}$ correspond to the decomposition of the line integrals in terms of degrees of freedom. In particular, Stokes’ formula (38) yields